# Deep Learning based Visually Rich Document Content Understanding: A Survey

### v1

@@ 6.1.2. Visual Representation
Visual information provides layout, structural insights, and rich contextual clues, making it easier for humans to interpret and prioritize content and resulting in a more comprehensive reading experience. Based on the methods used for encoding visual information, we categorize them into two main types: CNN-based and Vision Transformer-based approaches.
@@
CNN-based Vision Encoding methods involve first acquiring Region of Interest (RoI) bounding boxes and then applying RoI-pooling and RoI-Align on pretrained CNN backbones (e.g., Faster-RCNN or Mask-RCNN) to extract the region features. Many frameworks (Xu et al., 2020a; Li et al., 2021b; Luo et al., 2022; Ding et al., 2024a; Zhai et al., 2023; Ding et al., 2023b; Gu et al., 2021) utilize word or entity-level RoIs to effectively extract visual features of target regions and leverage the implicit knowledge embedded in pretrained backbones. However, acquiring the bounding boxes of words or entities incurs additional costs. Therefore, several frameworks (Xu et al., 2020b, 2021; Bai et al., 2022; Appalaraju et al., 2021) directly use image patch bounding RoIs to extract visual features and learn contextually with other modalities.
@@
After acquiring the visual features, they are typically fed into a transformer framework to fuse multimodal information, which can create significant computational bottlenecks. Additionally, acquiring high-quality RoIs of words or entities requires supervised training. To address these challenges, LayoutLMv3, inspired by ViT (Dosovitskiy et al., 2020), introduces a transformer-only framework. This approach applies a linear layer to project flattened patch pixels, which are then fed into a multimodal transformer to contextually learn with other modalities. This method reduces the number of parameters and simplifies the preprocessing steps, making it more efficient and adopted by many recent frameworks (Tang et al., 2022; Appalaraju et al., 2024; Mao et al., 2024). However, this encoding method cannot take advantage of implicit knowledge in pre-trained frameworks and typically requires extensive pre-trained.
6.1.3. Layout Representation
Figure 6. Commonly adopted layout information encoding approaches.
@@ 6.1.3. Layout Representation
2D positional encoding, first introduced by LayoutLM (Xu et al., 2020a), and widely adopted by many visually rich document understanding (VRDU) models, allows the model to be aware of the relative spatial positions within a document. In this approach, document elements are normalized and discretized into integer ranges, and two separate embedding layers are used to encode the x and y coordinates, respectively. Despite its widespread use in models like (Xu et al., 2020a, b, 2021; Huang et al., 2022; Tu et al., 2023; Wang et al., 2022b; Bai et al., 2022; Liao et al., 2023)this method encodes x and y coordinates individually, making it challenging to represent continuous 2D space and capture special correlations between document elements. Some models (Ding et al., 2023a; Wang et al., 2020b) that follow the approach of LXMERT (Tan and Bansal, 2019) utilize linear projection to update the x and y coordinates of the normalized bounding box coordinates simultaneously. To address the limitations of absolute positional encoding and incorporate relative positional correlations, other models introduce spatial-aware attention mechanisms (Hong et al., 2021; Appalaraju et al., 2021; Wang et al., 2022b). These mechanisms enable vanilla self-attention to learn spatial dependencies effectively.
6.2. Multi-modality Fusion
Figure 7. Commonly adopted multi-modality fusion methods.
@@ 6.2. Multi-modality Fusion
After acquiring multimodal representations, it is important to explore effective fusion methods to integrate textual, visual, and layout information. This integration improves document understanding and boosts performance on downstream tasks. The straightforward integration methods, as shown in Figure 7, include the additive and concatenation of the feature vectors. For example, additive integration sums layout information with corresponding textual or visual token representations (Xu et al., 2020a, b; Huang et al., 2022; Li et al., 2021b), while concatenation merges visual and textual features of document entities (e.g. tables) (Ding et al., 2024a, 2023b). However, these methods require one-to-one correlations and alternative approaches are needed when such correlations are not available. Consequently, self-attention and cross-attention mechanisms are widely adopted to enhance each modality by learning inter- and intramodality contexts. These mechanisms are commonly used in frameworks (Xu et al., 2020b; Huang et al., 2022; Tu et al., 2023; Bai et al., 2022; Xu et al., 2021) that integrate patch-level visual embeddings with textual features for contextual learning. Novel self-attention (Appalaraju et al., 2021; Wang et al., 2022b) and cross-attention (Gu et al., 2021; Li et al., 2021b; Zhai et al., 2023) methods have been proposed to fuse multimodal information more effectively. Apart from model-based fusion approaches, self-supervised and joint learning tasks are also effective for integrating multi-aspect features. Self-supervised pretraining tasks such as Masked Visual-Language Modeling (Wang et al., 2022b; Xu et al., 2020a, b; Li et al., 2021a; Huang et al., 2022; Zhai et al., 2023), Text-Image Alignment (Appalaraju et al., 2021), Text-Layout Pairing (Wang et al., 2022b), and Text-Image Matching (Xu et al., 2020b, 2021) can significantly enhance multimodal information fusion. These methods require large-scale pretraining to learn cross-modality semantic correlations. Joint learning methods, often used in OCR-free frameworks (Wang et al., 2021a; Kim et al., 2022; Davis et al., 2022; Yu et al., 2022), design auxiliary text detection or recognition tasks to fuse textual and visual information. This approach reduces pre-processing during inference and addresses mis-ordering sequence issues. However, these methods generally underperform compared to OCR-dependent models and involve additional training costs.
6.3. Model Architecture
6.3.1. Transformer in VRDU
@@
Referring to the models introduced in Section 3 and Section 3, transformers have become extensively utilized in VRDU tasks, attaining state-of-the-art performance due to several key advantages. Firstly, the attention mechanism effectively captures long-range dependencies within the multimodal information, including text, vision, and layout. Furthermore, the inherent scalability of transformers improves self-supervised learning on large-scale datasets (e.g. IIT-CDIP (Lewis et al., 2006)), allowing them to handle diverse types and formats of documents, capturing more intricate and comprehensive features of the documents. Based on the transformer architecture used, VRDU models can be divided into two categories: encoder-only and encoder-decoder-based models. The first encoder-only model, LayoutLM (Xu et al., 2020a), was inspired by BERT (Devlin, 2018) and uses various pretraining tasks to allow the bidirectional transformer encoder to capture more textual and layout information. Following LayoutLM, more pretrained VRDU models with encoder-only, layout-aware (Wang et al., 2022b; Hong et al., 2021; Chen et al., 2022; Wang et al., 2022b; Tu et al., 2023; Li et al., 2021a) or visual integrated (Xu et al., 2020b, 2021; Appalaraju et al., 2021; Huang et al., 2022) have been proposed. These models are pretrained on various tasks to enhance their understanding of document structures. Encoder-only models demonstrate remarkable performance in sequence tagging and document classification tasks. However, they face challenges related to heavy annotation requirements and low readability, and they struggle with generative tasks such as abstractive question answering. Additionally, OCR errors can complicate the extraction of accurate information from input text. Furthermore, the fixed maximum input length of encoder-only frameworks limits their ability to handle long document inputs effectively. To overcome the limitations of encoder-only frameworks in generative QA and KIE, several encoder-decoder models (Powalski et al., 2021; Tang et al., 2022; Appalaraju et al., 2024; Mao et al., 2024) have been developed, but they still depend on costly OCR tools which can introduce errors affecting performance. OCR-free frameworks (Kim et al., 2022; Davis et al., 2022; Cao et al., 2023a; Yu et al., 2022) address this issue by using vision encoders and text decoders for end-to-end processing. For long documents, T5-based encoder-decoder models (Raffel et al., 2020) have been proposed to effectively handle multipage contexts.
6.3.2. CNNs in VRDU
@@
In VRDU, CNNs are employed as the core framework for extracting feature maps from document images (Yang et al., 2017; Palm et al., 2019) and character grids (Katti et al., 2018), benefiting from their strong local feature extraction capabilities. Joint-learning frameworks (Wang et al., 2021a, a) use CNNs as a backbone to integrate OCR and KIE tasks, combining visual and textual information through auxiliary tasks. Some OCR-free pretrained frameworks (Davis et al., 2022; Yu et al., 2022) also utilize CNNs as vision encoders to extract visual feature maps. However, CNNs struggle to capture long-range dependencies due to their localized receptive fields. To address this, StrucTexTv2 (Yu et al., 2022) combines CNN-extracted feature maps with a transformer to capture global contextual information. Additionally, CNNs are commonly used to extract visual features from regions of interest (RoI) with RoI Align(Xu et al., 2020a, b, 2021; Huang et al., 2022; Tu et al., 2023; Wang et al., 2022b; Bai et al., 2022; Liao et al., 2023). Although CNNs can accurately capture region-specific visual cues and leverage pre-trained knowledge from general domains, they require extra processing to obtain RoI bounding boxes, unlike vision transformers (Huang et al., 2022; Tang et al., 2022; Appalaraju et al., 2024; Mao et al., 2024), which operate directly on patch pixel values.
6.3.3. Graphs in VRDU
@@
VRDs feature complex spatial and logical structures. The spatial structure shows the layout and positional relationships, such as a Title above a Paragraph and a Caption near a Figure or Table. The logical structure denotes semantic and hierarchical connections, like a Title being the parent of a Paragraph and a Caption describing a related Table or Figure. Graph-based frameworks explicitly encode these relationships using node and edge representations; therefore, Graph Neural Networks (GNNs) are widely used in VRDU models (Lee et al., 2022b; Shi et al., 2023; Luo et al., 2022; Zhang et al., 2021) to encode the spatial and logical representations. Although GNNs effectively capture domain-specific knowledge, they struggle with scalability and general domain knowledge pretraining. To address this, some frameworks (Li et al., 2023; Zhang et al., 2022) use attention masks or biases to mimic relationships between document elements, blending attention mechanisms with explicit relational modelling.
6.4. Pretraining Mechanisms
@@
By performing various pretraining tasks, a model can enhance its generalization ability through extensive datasets and prior training. Inspired by advances in pre-training language (Devlin, 2018; Brown et al., 2020) and vision models (He et al., 2016; Dosovitskiy et al., 2020), numerous pretraining tasks for VRDU have been developed that are typically trained in large-scale document collections. This section will summarize the pretraining techniques and datasets commonly used for VRDU pretraining.
6.4.1. Pretraining Tasks
@@
Masked Information Modelling (MIM) is first introduced by Masked Language Modelling in BERT (Devlin, 2018), which randomly masks 15% workpiece tokens and requires the model to predict the masked tokens. Some models use multi-source inputs, such as XDoc (Chen et al., 2022) and MarkupLM (Li et al., 2022a), which directly adopts Masked Language Modeling as a pretraining task on plain text or markdown text subsets. Some methods optimise MLM by changing masking wordpieces to whole words (Whole Word Masking by LayoutMask (Tu et al., 2023)) or all tokens belong to one randomly generated text block, named Area-Masked Language Modelling (Hong et al., 2021). Similar strategies can also be applied to mask entity-level textual representation, e.g. Masked Sentence Modelling (Gu et al., 2021). Those language-targeted masked information modelling to improve the language understanding ability of VRDU models. Except for language-focused masking strategies, vision (Gu et al., 2021; Yu et al., 2022) and layout-focused (Tu et al., 2023; Wang et al., 2022b) strategies are also adopted. Moreover, masked information modelling is an effective method to boost cross-modality understanding. LayoutLM (Xu et al., 2020a) introduces a Masked Visual-Language Modelling which allows using kept visual/layout information and contextual text content to predict the masked word-pieces, adopted by many VRDU pretrained models (Wang et al., 2022b; Xu et al., 2020b; Li et al., 2021a; Huang et al., 2022; Zhai et al., 2023; Appalaraju et al., 2021). Similarly, some visual token masked models leverage multimodal information to reconstruct the view tokens, e.g. Learn to Reconstruct (Appalaraju et al., 2021), Masked Image Modeling (Huang et al., 2022). Additionally, some models mask multimodal features simultaneously to conduct a cross-modality masking (Li et al., 2021b; Liao et al., 2023). Masking Information Modelling may have limitations on bias in masking strategies, thus some frameworks (Tu et al., 2023) may try different masking ratios or strategies to improve the training effectiveness. Additionally, other common concerns about MIM also include training efficiency and lack of structural information. Thus, other pretraining methods are introduced to mitigate the limitations.
@@
Cross-modality Aligning Although some Masked Information Modelling methods could effectively boost the cross-modality understanding, implicit contextual learning is limited to capturing explicit alignment between different modalities. Thus, few cross-modality aligning methods are introduced to enhance the modality interaction. To enhance the text-image interactive learning, Text-Image Alignment is adopted (Xu et al., 2020b, 2021) which covers the image region of token lines to predict whether the image region of the target token line is covered or not. LayoutLMv3 (Huang et al., 2022) expands from covering the image region only to image/text tokens to further enhance interactive learning. Vision-language alignment (Gu et al., 2021) and Text-Image Matching (Xu et al., 2020b, 2021) target to predict whether image-text features belong to the same region or not. DocFormer (Appalaraju et al., 2021) tends to predict the text content of the paired images.
@@
Other Pretraining Techniques: Some pretraining tasks are introduced to further enhance the understanding of specific modalities. To enhance layout information understanding, StructuralLM (Li et al., 2021a) and WOKONG-READER (Bai et al., 2022) introduce Cell Position Classification and Textline Grid Matching to predict which located grids are of each cell or textline. Fast-StrucText (Zhai et al., 2023) introduces a graph-based token relation method to predict a spatial correlation between token pairs. MarkupLM (Li et al., 2022a) leverage the benefits from markup files to predict the logical parent-child relation between nodes by introducing a Node Relation Prediction. Contrastive learning based strategies are adopted to conduct single (Gu et al., 2021) or cross-modality contrastive (Textline-Region Contrastive Learning) (Bai et al., 2022) learning.
6.4.2. Pretraining Datasets

### v2

@@ 6.1.2 Visual Representation
Visual information provides layout, structural insights, and rich contextual clues, making it easier for humans to interpret and prioritize content and resulting in a more comprehensive reading experience. Based on the methods used for encoding visual information, we categorize them into two main types: CNN-based and Vision Transformer-based approaches, as represented by Figure 5.
@@
CNN-based Vision Encoding. Methods involve first acquiring Region of Interest (RoI) bounding boxes and then applying RoI-pooling and RoI-Align on pretrained CNN backbones (e.g., Faster-RCNN or Mask-RCNN) to extract the region features. Many frameworks [layoutlm, selfdoc, docgcn, mmvqa, faststructext, pdfvqa, udoc] utilize word or entity-level RoIs to effectively extract visual features of target regions and leverage the implicit knowledge embedded in pretrained backbones. However, acquiring the bounding boxes of words or entities incurs additional costs. Therefore, several frameworks [layoutlmv2, layoutxlm, wukong, docformer] directly use image patch bounding RoIs to extract visual features and learn contextually with other modalities.
@@
Tranformer-based Vision Encoding. After acquiring the visual features, they are typically fed into a transformer framework to fuse multimodal information, which can create significant computational bottlenecks. Additionally, acquiring high-quality RoIs of words or entities requires supervised training. To address these challenges, LayoutLMv3, inspired by ViT [vit], introduces a transformer-only framework. This approach applies a linear layer to project flattened patch pixels, which are then fed into a multimodal transformer to contextually learn with other modalities. This method reduces the number of parameters and simplifies the preprocessing steps, making it more efficient and adopted by many recent frameworks [udop, docformerv2, vitlp]. However, this encoding method cannot take advantage of implicit knowledge in pre-trained frameworks and typically requires extensive pre-trained.
6.1.3 Layout Representation
Figure 6: Overview of layout encoding methods. Left: Sinusoidal positional encoding applied to bounding box coordinates $\text{PE}(bbox_{E})=[\text{PE}(x_{1});\text{PE}(y_{1});\text{PE}(x_{2});\text{PE}(y_{2})]\in\mathbb{R}^{d}$. Middle: Linear projection of layout features $P_{E}=\text{Linear}(bbox_{E})\in\mathbb{R}^{d}$. Right: Attention bias, where layout-aware attention is modified via $\alpha(E_{1},E_{2})=\alpha+f(bbox_{E_{1}},bbox_{E_{2}})$.
@@ 6.1.3 Layout Representation
2D positional encoding, first introduced by LayoutLM [layoutlm], and widely adopted by many VRD-CU models, allows the model to be aware of the relative spatial positions within a document. In this approach, document elements are normalized and discretized into integer ranges, and two separate embedding layers are used to encode the x and y coordinates, respectively. Despite its widespread use in models like [layoutlm, layoutlmv2, layoutxlm, layoutlmv3, layoutmask, lilt, wukong, doctr]this method encodes x and y coordinates individually, making it challenging to represent continuous 2D space and capture special correlations between document elements. Some models [formnlu, docstruct] that follow the approach of LXMERT [lxmert] utilize linear projection to update the x and y coordinates of the normalized bounding box coordinates simultaneously. To address the limitations of absolute positional encoding and incorporate relative positional correlations, other models introduce spatial-aware attention mechanisms [bros, docformer, lilt]. These mechanisms enable vanilla self-attention to learn spatial dependencies effectively.
6.2 Multi-modality Fusion
Figure 7: Overview of multimodal fusion strategies. Left: Additive ($e_{1}=t_{1}+v_{1}$) and concatenation ($e_{1}=t_{1}\oplus v_{1}$) fusion, with $t_{1},v_{1}\in\mathbb{R}^{d}$. Middle: Pretraining via $\mathcal{F}(v,p)\rightarrow t$ or joint-learning by minimizing $\mathcal{L}_{\text{task1}}+\mathcal{L}_{\text{task2}}$. Right: Self-attention and cross-attention mechanisms using $\text{Attention}(Q,K,V)=\text{softmax}\left(QK^{\top}/{\sqrt{d_{k}}}\right)V$.
@@ 6.2 Multi-modality Fusion
After acquiring multimodal representations, it is important to explore effective fusion methods to integrate textual, visual, and layout information. This integration improves document understanding and boosts performance on downstream tasks. The straightforward integration methods, as shown in Figure 7, include the additive and concatenation of the feature vectors. For example, additive integration sums layout information with corresponding textual or visual token representations [layoutlm, layoutlmv2, layoutlmv3, selfdoc], while concatenation merges visual and textual features of document entities (e.g. tables) [mmvqa, pdfvqa]. However, these methods require one-to-one correlations and alternative approaches are needed when such correlations are not available. Consequently, self-attention and cross-attention mechanisms are widely adopted to enhance each modality by learning inter- and intramodality contexts. These mechanisms are commonly used in frameworks [layoutlmv2, layoutlmv3, layoutmask, wukong, layoutxlm] that integrate patch-level visual embeddings with textual features for contextual learning. Novel self-attention [docformer, lilt] and cross-attention [udoc, selfdoc, faststructext] methods have been proposed to fuse multimodal information more effectively. Apart from model-based fusion approaches, self-supervised and joint learning tasks are also effective for integrating multi-aspect features. Self-supervised pretraining tasks such as Masked Visual-Language Modeling [lilt, layoutlm, layoutlmv2, structurallm, layoutlmv3, faststructext], Text-Image Alignment [docformer], Text-Layout Pairing [lilt], and Text-Image Matching [layoutlmv2, layoutxlm] can significantly enhance multimodal information fusion. These methods require large-scale pretraining to learn cross-modality semantic correlations. Joint learning methods, often used in OCR-free frameworks [ephoie, donut, dessurt, structextv2], design auxiliary text detection or recognition tasks to fuse textual and visual information. This approach reduces pre-processing during inference and addresses mis-ordering sequence issues. However, these methods generally underperform compared to OCR-dependent models and involve additional training costs.
6.3 Model Architecture
6.3.1 Transformer in VRD-CU
@@
Referring to the models introduced in Section 3 and Section 3, transformers have become extensively utilized in VRD-CU tasks, attaining state-of-the-art performance due to several key advantages. Firstly, the attention mechanism effectively captures long-range dependencies within the multimodal information, including text, vision, and layout. Furthermore, the inherent scalability of transformers improves self-supervised learning on large-scale datasets (e.g. IIT-CDIP [iitcdip]), allowing them to handle diverse types and formats of documents, capturing more intricate and comprehensive features of the documents. Based on the transformer architecture used, VRD-CU models can be divided into two categories: encoder-only and encoder-decoder-based models. The first encoder-only model, LayoutLM [layoutlm], was inspired by BERT [bert] and uses various pretraining tasks to allow the bidirectional transformer encoder to capture more textual and layout information. Following LayoutLM, more pretrained VRD-CU models with encoder-only, layout-aware [lilt, bros, xdoc, lilt, layoutmask, structurallm] or visual integrated [layoutlmv2, layoutxlm, docformer, layoutlmv3] have been proposed. These models are pretrained on various tasks to enhance their understanding of document structures. Encoder-only models demonstrate remarkable performance in sequence tagging and document classification tasks. However, they face challenges related to heavy annotation requirements and low readability, and they struggle with generative tasks such as abstractive question answering. Additionally, OCR errors can complicate the extraction of accurate information from input text. Furthermore, the fixed maximum input length of encoder-only frameworks limits their ability to handle long document inputs effectively. To overcome the limitations of encoder-only frameworks in generative QA and KIE, several encoder-decoder models [tilt, udop, docformerv2, vitlp] have been developed, but they still depend on costly OCR tools which can introduce errors affecting performance. OCR-free frameworks [donut, dessurt, serum, structextv2] address this issue by using vision encoders and text decoders for end-to-end processing. For long documents, T5-based encoder-decoder models [t5] have been proposed to effectively handle multipage contexts.
6.3.2 CNNs in VRD-CU
@@
In VRD-CU, CNNs are employed as the core framework for extracting feature maps from document images [yang2017learning, palm2019attend] and character grids [chargrid], benefiting from their strong local feature extraction capabilities. Joint-learning frameworks [ephoie, ephoie] use CNNs as a backbone to integrate OCR and KIE tasks, combining visual and textual information through auxiliary tasks. Some OCR-free pretrained frameworks [dessurt, structextv2] also utilize CNNs as vision encoders to extract visual feature maps. However, CNNs struggle to capture long-range dependencies due to their localized receptive fields. To address this, StrucTexTv2 [structextv2] combines CNN-extracted feature maps with a transformer to capture global contextual information. Additionally, CNNs are commonly used to extract visual features from regions of interest (RoI) with RoI Align[layoutlm, layoutlmv2, layoutxlm, layoutlmv3, layoutmask, lilt, wukong, doctr]. Although CNNs can accurately capture region-specific visual cues and leverage pre-trained knowledge from general domains, they require extra processing to obtain RoI bounding boxes, unlike vision transformers [layoutlmv3, udop, docformerv2, vitlp], which operate directly on patch pixel values.
6.3.3 Graphs in VRD-CU
@@
VRDs feature complex spatial and logical structures. The spatial structure shows the layout and positional relationships, such as a Title above a Paragraph and a Caption near a Figure or Table. The logical structure denotes semantic and hierarchical connections, like a Title being the parent of a Paragraph and a Caption describing a related Table or Figure. Graph-based frameworks explicitly encode these relationships using node and edge representations; therefore, GNNs are widely used in VRD-CU models [formnet, layoutgcn, docgcn, zhang2021entity] to encode the spatial and logical representations. Although GNNs effectively capture domain-specific knowledge, they struggle with scalability and general domain knowledge pretraining. To address this, some frameworks [graphlayoutlm, graphdoc] use attention masks or biases to mimic relationships between document elements, blending attention mechanisms with explicit relational modelling.
6.4 Pretraining Mechanisms
@@
By performing various pretraining tasks, a model can enhance its generalization ability through extensive datasets and prior training. Inspired by advances in pre-training language [bert, gpt3] and vision models [resnet, vit], numerous pretraining tasks for VRD-CU have been developed that are typically trained in large-scale document collections. This section will summarize the pretraining techniques and datasets commonly used for VRD-CU pretraining.
6.4.1 Pretraining Tasks
@@
Masked Information Modelling (MIM) is first introduced by Masked Language Modelling in BERT [bert], which randomly masks 15% workpiece tokens and requires the model to predict the masked tokens. Some models use multi-source inputs, such as XDoc [xdoc] and MarkupLM [markuplm], which directly adopts Masked Language Modeling as a pretraining task on plain text or markdown text subsets. Some methods optimise MLM by changing masking wordpieces to whole words (Whole Word Masking by LayoutMask [layoutmask]) or all tokens belong to one randomly generated text block, named Area-Masked Language Modelling [bros]. Similar strategies can also be applied to mask entity-level textual representation, e.g. Masked Sentence Modelling [udoc]. Those language-targeted masked information modelling to improve the language understanding ability of VRD-CU models. Except for language-focused masking strategies, vision [udoc, structextv2] and layout-focused [layoutmask, lilt] strategies are also adopted. Moreover, masked information modelling is an effective method to boost cross-modality understanding. LayoutLM [layoutlm] introduces a Masked Visual-Language Modelling which allows using kept visual/layout information and contextual text content to predict the masked word-pieces, adopted by many VRD-CU pretrained models [lilt, layoutlmv2, structurallm, layoutlmv3, faststructext, docformer]. Similarly, some visual token masked models leverage multimodal information to reconstruct the view tokens, e.g. Learn to Reconstruct [docformer], Masked Image Modeling [layoutlmv3]. Additionally, some models mask multimodal features simultaneously to conduct a cross-modality masking [selfdoc, doctr]. Masking Information Modelling may have limitations on bias in masking strategies, thus some frameworks [layoutmask] may try different masking ratios or strategies to improve the training effectiveness. Additionally, other common concerns about MIM also include training efficiency and lack of structural information. Thus, other pretraining methods are introduced to mitigate the limitations.
@@
Cross-modality Aligning Although some Masked Information Modelling methods could effectively boost the cross-modality understanding, implicit contextual learning is limited to capturing explicit alignment between different modalities. Thus, few cross-modality aligning methods are introduced to enhance the modality interaction. To enhance the text-image interactive learning, Text-Image Alignment is adopted [layoutlmv2, layoutxlm] which covers the image region of token lines to predict whether the image region of the target token line is covered or not. LayoutLMv3 [layoutlmv3] expands from covering the image region only to image/text tokens to further enhance interactive learning. Vision-language alignment [udoc] and Text-Image Matching [layoutlmv2, layoutxlm] target to predict whether image-text features belong to the same region or not. DocFormer [docformer] tends to predict the text content of the paired images.
@@
Other Pretraining Techniques: Some pretraining tasks are introduced to further enhance the understanding of specific modalities. To enhance layout information understanding, StructuralLM [structurallm] and WOKONG-READER [wukong] introduce Cell Position Classification and Textline Grid Matching to predict which located grids are of each cell or textline. Fast-StrucText [faststructext] introduces a graph-based token relation method to predict a spatial correlation between token pairs. MarkupLM [markuplm] leverage the benefits from markup files to predict the logical parent-child relation between nodes by introducing a Node Relation Prediction. Contrastive learning-based strategies are adopted to conduct single [udoc] or cross-modality contrastive (Textline-Region Contrastive Learning) [wukong] learning.
6.4.2 Pretraining Datasets